OpenStudy (anonymous):

Sue is twice as old as Sam, six years ago Sue was three times as old as Sam. How old are Sue and Sam now?

OpenStudy (cggurumanjunath):

let sam's age =x sue age =2x

OpenStudy (cggurumanjunath):

is it clear @AnImEfReaK

OpenStudy (cggurumanjunath):

?

OpenStudy (anonymous):

I found that

OpenStudy (anonymous):

But I can't find another equation to put on the other side.

OpenStudy (anonymous):

The last part of the qn confuses me '' six years ago Sue was three times as old as Sam''

OpenStudy (cggurumanjunath):

six yrs ago sue age =3sam ; six years ago Sue was three times as old as Sam 2x-6=3x-6

OpenStudy (cggurumanjunath):

ok ?

OpenStudy (anonymous):

I don't get it...

OpenStudy (cggurumanjunath):

2x-6=3(x-6)

OpenStudy (cggurumanjunath):

sorry

OpenStudy (anonymous):

Why is it 2x-6?

OpenStudy (anonymous):

I get what you did at the end now.

OpenStudy (cggurumanjunath):

2x-6 b'coz ,6yrs ago which is past so 6yrs must be subtracted.

OpenStudy (cggurumanjunath):

clear ?

OpenStudy (anonymous):

I got 2x = 3(x-6).

OpenStudy (cggurumanjunath):

2x-6 not 2x in lhs !

OpenStudy (cggurumanjunath):

u have to subtract age on both sides !

OpenStudy (anonymous):

But Sue was not twice as old as Sam 6 years ago.

OpenStudy (anonymous):

It's so confusing...

OpenStudy (cggurumanjunath):

age at 6 yrs back means u have to subtract 6 yrs on both sides .

OpenStudy (anonymous):

Okay then it applies to both.

OpenStudy (cggurumanjunath):

yes

OpenStudy (anonymous):

Thanks @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

wc.